40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 30 17:01:08 2011, Larisa Thorne, Update, VIDEO, Cable laying...continued 
    Reply  Thu Mar 31 08:31:44 2011, Suresh, Update, VIDEO, Cable laying...continued P1070489.JPGP1070494.JPG
Message ID: 4474     Entry time: Thu Mar 31 08:31:44 2011     In reply to: 4462
Author: Suresh 
Type: Update 
Category: VIDEO 
Subject: Cable laying...continued 

The video work has crossed a milestone.    

Kiwamu and Steve have shifted the three quads from the control room to the Video MUX rack (1Y1) and have wired them to the MUX.

The monitors in the control room have been repositioned and renumbered.  They are now connected directly to the MUX. 

Please see the new cable list for the input and output channels on the MUX.

As of today, all cables according the new plan are in place.  Their status   indicated on the wiki page above is not verified .  Please ignore that column for now, we will be updating that soon.

I shifted the MC1F/MC3F camera and the MC2F cameras onto the new cables.  Also connected the monitors at the BS chamber and end of the X arm to their respective cables.  I have removed the RG58B BNC (black) cables running from MC2 to BS and from ETMXF to the top of the Flow Bench. 

Some of the old video cables are still in place but are not used.   We might consider removing them to clear up the clutter. 

Some of the video cables in use are orange and if the lab's  cable color code is to be enforced these will have to be replaced with blue ones..

Some of the cables in use running from the MUX to the monitor in the control room are the white 50 Ohm variety.  There are also black RG59 Cables running the same way ( we have surplus cables in that path)  and we have to use those instead of the white ones. 

There are a number of tasks remaining:

a)  The inputs from the various existing cameras have to be verified. 

b) There are quite a few cameras which are yet to be installed.

c) The Outputs may not not be connected to their monitors.  That the monitors may still be connected to an old cable which is not connected to the MUX.  The new cable should be lying around close by.  So if you see a blank monitor please connect it to its new cable. 

d) The status column on the wiki page has to be updated.

e) Some of the currently in place may need to be replaced and some need to be removed.  We need to discuss our priorities and come up with a plan for that.

After checking everything we can certify that the video cabling system is complete.

I would like Joon Ho to take care of this verification+documenting process and declaring that the job is complete. 

 

Steve attached these two pictures.

Attachment 1: P1070489.JPG  2.934 MB  Uploaded Thu Mar 31 11:56:33 2011  | Hide | Hide all
P1070489.JPG
Attachment 2: P1070494.JPG  2.906 MB  Uploaded Thu Mar 31 11:57:23 2011  | Hide | Hide all
P1070494.JPG
ELOG V3.1.3-